4) A 40.0 lb, 10.0 ladder leans against a smooth vertical wall as shown. It makes an angle of 60.0° with the horizontal floorplease solve all
Answer #1

4 Draw the free face Body diagram 40 lb NA 180 lb x let, t be the distance from from point B before sliding begins NS Vertica

Take moment about B, EMB=0 to find distance a the 10 Cos 60 [Cow [MAX 10 Sin 68] - [HQ* (10 - [180] [2006] 8.66 NA 100.-90.xF and the weight, W= ng 150 x 9.81 - 14T15N cose-la sliding up of mass 147165 SA Find the prictional farce, FA= Ms NA uthor =Resolve along horizontal directan 2Fx=0 + F NASIA 3o. - FA cos30=0 +FA = 0.5NA -0.303 NA = 0 11 F = 0.803 NA F= 0.803X 2129.4-14715 + NA COS 38 + FA sin 30 =0 - 149 15+ NA (8838 +6:35 NA Sin 30=0 NA = 1413-51 = 1413.5 IN Horizontal Resove along weutiNow, Elementaal Area dA = drody = (29-4²) dy Find the centroid of the Element gła : 29+y [yty 2 2 Find the centroid of the to

the Find the centroid of Ica- Find Elemental Area dA = dredy = (29-6?). Find the centroid y Element ☺ = y. Find the centroid# Find the centroid ý distace given Equation, 2 ay Z= 3 ry 7 / 3 = Jy > z=ay Elemental volume, dv = (1 z) dy qy) dy Find the

I have provided conceptually correct answer.

If you have any doubts please ask me.

Please provide positive rating and thank you.